### v2

# uncompress

Abstract.
In this paper we investigate the degrees of irrationality of degenerations of $\epsilon$-lc Fano varieties of arbitrary dimensions. We show that given a generically $\epsilon$-lc klt Fano fibration $X\to Z$ of dimension $d$ over a smooth curve $Z$ such that $(X,tF)$ is lc for a positive real number $t$ where $F$ is the reduction of an irreducible central fibre of $X$ over a closed point $z\in Z$, then $F$ admits a rational dominant map $\pi\colon F\dashrightarrow C$ to a smooth projective variety $C$ with bounded degree of irrationality depending only on $\epsilon,d,t$ such that the general fibres of $\pi$ are irreducible and rational. This proves the generically bounded case of a conjecture proposed by the first author and Loginov for log Fano fibrations of dimensions greater than three. One of the key ingredients in our proof is to modify the generically $\epsilon$-lc klt Fano fibration $X\to Z$ to a toroidal morphism of toroidal embeddings with bounded general fibres.
2020 MSC: 14B05, 14D06, 14E30, 14J45, 14M20, 14M25.
1. Introduction
We work over an algebraically closed field $\mathbb{K}$ of characteristic zero unless stated otherwise.
Bounded irrationality of divisors on Fano fibrations. Given a family of varieties parameterised by a smooth curve, degenerations are the limit varieties of this family of varieties. Many geometric properties of varieties are preserved by passing to degenerations. For example, irreducible degenerations of uniruled varieties and rationally chain connected varieties are also uniruled and rationally chain connected; see [Kol96, Theorem IV.1.8, Theorem IV.3.5.3]. However, it is well-known that rational varieties can degenerate to non-rational varieties even for rational del Pezzo surfaces; for instance, a smooth cubic surface can degenerate to a cone over an elliptic curve which is a non-rational and singular del Pezzo surface. This phenomenon is closely related to the fact that the degeneration has log canonical (lc) but not Kawamata log terminal (klt) singularities.
In the paper [BL21], the first named author and Loginov study the non-rationality property of degenerations of klt del Pezzo surfaces and prove how far the components of the degenerations can be from being rational. More precisely, the boundedness of irrationality of degenerations is proved in [BL21] as follows. In particular, by [BL21, Theorem 1.1], the irrationality of degenerations of klt del Pezzo surfaces are bounded in a certain way even though the degenerations do not belong to a bounded family of surfaces, that is, the collection of degenerations can not be parameterised by a finite type scheme; see §2.9. Recall that a contraction $f\colon X\to Z$ of normal varieties is called a klt Fano fibration if $X$ has klt singularities and $K_{X}$ is ample over $Z$; see §2.5.
Theorem 1.1 ([BL21, Theorem 1.1]).
Fix a positive real number $t$. Assume that $f\colon X\to Z$ is a klt Fano fibration where $\dim X=3$ and $\dim Z=1$. Assume that $F$ is the reduction of an irreducible fibre and that $(X,tF)$ is lc. Then:
  • (i)
    $F$ is birational to $\mathbb{P}^{1}\times C$, where $C$ is a smooth projective curve with gonality $\operatorname{gon}(C)$ bounded depending only on $t$,
  • (ii)
    if $t>\frac{1}{2}$, then the genus $g(C)$ is bounded, and
  • (iii)
    if $t=1$, then the genus $g(C)\leq 1$.
Note that the genus $g(C)$ can be arbitrarily large if $t\leq\frac{1}{2}$; see [BL21, Example 2.3]. Moreover, Theorem 1.1 follows from a more general result for Fano type log Calabi-Yau fibrations where the base $Z$ of the fibration can have dimension $\geq 1$; see [BL21, Theorem 1.3]. Recall that the gonality $\operatorname{gon}(C)$ of a smooth curve $C$ is the smallest degree of the dominant morphisms of curves $C\to\mathbb{P}^{1}$. In higher dimensions, the degree of irrationality $\operatorname{irr}(D)$ of a variety $D$ is the least possible degree of dominant rational maps $D\dashrightarrow\mathbb{P}^{\dim D}$; see §2.16. The first named author and Loginov ask in [BL21] whether a similar result to Theorem 1.1 for klt Fano fibrations of higher dimensions should also hold.
Question 1.2 (cf. [BL21, Question 1.4]).
Fix a positive real number $t>0$ and natural number $d$. Suppose that $f\colon X\to Z$ is a klt Fano fibration over a smooth curve $Z$, where $\dim X=d$. Assume that $D$ is the reduction of an irreducible fibre of $f$ such that $(X,tD)$ is lc. Is it true that there is a rational map $D\dashrightarrow C$, where the general fibres are rationally connected and $C$ is a smooth projective variety with bounded degree of irrationality?
In this paper, we apply a very different approach from [BL21] (cf. [BL21, §5, §6]) via toroidal geometry to prove the boundedness of degrees of irrationality for degenerations of klt Fano fibrations of arbitrary dimensions with $\epsilon$-lc general fibres.
Theorem 1.3.
Fix positive real numbers $\epsilon,t>0$ and a natural number $d$. Assume that $f\colon X\to Z$ is a klt Fano fibration with $\dim X=d$ such that
  • (1)
    $Z$ is a smooth curve,
  • (2)
    $X$ is $\epsilon$-lc over the generic point of $Z$, and
  • (3)
    $F$ is the reduction of an irreducible fibre of $f$ and $(X,tF)$ is lc.
Then there is a dominant rational map $F\dashrightarrow C$ whose general fibres are irreducible and rational, and $C$ is a bounded smooth projective variety hence with bounded degree of irrationality.
Note that the general fibres of $F\dashrightarrow C$ in Theorem 1.3 are rational, which is a much stronger geometric property than the rational connectivity predicted in Question 1.2. Example 2.1 of [BL21] shows that Theorem 1.3 does not hold without assuming that $(X,tF)$ is lc. A similar result can be formulated when the fibre containing $F$ is not irreducible by using Fano type log Calabi-Yau fibrations; cf. [BL21, Theorem 1.3, Question 1.4].
As a corollary of Theorem 1.3, we show that every irreducible component of fibres of $\epsilon$-lc Fano fibrations admits the structure as in Theorem 1.3. Note that in this case, we do not prescribe a positive real number $t>0$ so that $(X,tF)$ is lc.
Theorem 1.4.
Let $\epsilon$ be a positive real number, and let $d$ be a natural number. Let $f\colon X\to Z$ be a Fano fibration with $\dim X=d$, where $Z$ is a smooth curve, and $X$ is $\epsilon$-lc. Let $F$ be an irreducible component of a closed fibre of $f$. Then there is a dominant rational map $F\dashrightarrow C$ whose general fibres are irreducible and rational, and $C$ is a bounded smooth projective variety hence with bounded degree of irrationality.
Relative toroidalisation and sketch of the proof. One of the key ingredients in our proof of Theorem 1.3 is that we can modify the family $f\colon X\to Z$ to a generically bounded toroidal morphism of strict toroidal embeddings. A general fibre of $X\to Z$ in Theorem 1.3 is an $\epsilon$-lc Fano variety which is bounded by [Bir21, Theorem 1.1]. Thus, there is a birational map $\phi\colon X\dashrightarrow Y$ over $Z$ such that every fibre of $g\colon Y\to Z$ is bounded, that is, $Y\to Z$ is relatively bounded; see §2.13. Moreover, $\phi$ can be chosen so that it does not contract any curve over the generic point of $Z$ as general fibres of $f$ are already bounded. In §4, we show that there exists a toroidal morphism $f^{\prime}$ fitting into the commutative diagram
[…]
where the left vertical arrow $f^{\prime}$ is a toroidal morphism of strict toroidal embeddings (see §3.3 and §3.6), $m_{X}$ and $m_{Z}$ are projective birational morphisms, and the general fibres of $f^{\prime}$ are bounded.
Note that the toroidal modification exists for any projective surjective morphism of varieties by [AK00, ADK13]; see Theorem 3.10. However, here we require a stronger condition on the boundedness of general fibres of $f^{\prime}$, which will be deduced from Theorem 3.12. Similar constructions are applied in [Bir23a] where $m_{X}$ may not be birational. The techniques related to toroidal embeddings are also used in [Jia23]; see, for example, [Jia23, §3].
Denote by $C^{\prime}$ the centre of $F$ on $Y^{\prime}$. Decreasing $t$ if necessary, we can assume that $t$ is a positive rational number. By taking an $n$-complement of $K_{X}+tF$ for some $n\in\mathbb{N}$ depending only on $t$ (see §2.6), it will be shown in §4 that the toroidal modification $f^{\prime}$ can be taken so that these additional properties are also satisfied:
  • (i)
    the support of the fibre of $f^{\prime}$ over $z\in Z\cong Z^{\prime}$ is contained in the toroidal boundary $D^{\prime}\coloneqq Y^{\prime}\setminus U_{Y^{\prime}}$, in particular, $C^{\prime}$ is contained in $D^{\prime}$, and
  • (ii)
    the log discrepancy $a(F,Y^{\prime},D^{\prime})$ is equal to zero.
[…]
Let $C$ be a bounded resolution of $C^{\prime}$, then there is a rational map $\pi\colon F\dashrightarrow C$. Now, $C$ is a smooth projective variety with bounded degree of irrationality since it is bounded; cf. Lemma 2.17. Moreover, as $C^{\prime}$ is an lc centre of the strict toroidal couple $(Y^{\prime},D^{\prime})$, a general fibre of $\pi$ is irreducible and rational; see Proposition 3.5. This concludes Theorem 1.3.
The following result compares Theorem 1.3 and [BL21, Theorem 1.1] when $\dim X=3$.
Corollary 1.5 (cf. [BL21, Theorem 1.1]).
Fix positive real numbers $\epsilon,t$. Let $f\colon X\to Z$ be a klt Fano fibration where $\dim X=3$ and $\dim Z=1$ such that $X$ is $\epsilon$-lc over the generic point of $Z$. Let $F$ be the reduction of an irreducible fibre of $f$ such that $(X,tF)$ is lc. Then $F$ is birational to $\mathbb{P}^{1}\times E$, where $E$ is a smooth projective curve with gonality $\operatorname{gon}(E)$ and genus $g(E)$ bounded depending only on $\epsilon$ and $t$.
Proof.
From the discussion as above, there is a rational map $\pi\colon F\dashrightarrow C$ such that
  • (i)
    $C$ is a bounded smooth projective variety, and
  • (ii)
    a general fibre of $\pi$ is irreducible and rational.
Then the dimension of $C$ gives several possibilities on the structure of $F$:
  • (1)
    if $\dim C=0$, then $F$ is a rational surface,
  • (2)
    if $\dim C=1$, then $C$ has bounded gonality $\operatorname{gon}(C)$ and genus $g(C)$, and a general fibre of $\pi$ is isomorphic to $\mathbb{P}^{1}$, hence $F$ is birational to $\mathbb{P}^{1}\times C$, and
  • (3)
    if $\dim C=2$, then $F$ is birationally bounded.
Note that $F$ must be ruled as it is a degeneration of a rational surface (see [Kol96, Theorem IV.1.6]), so $F$ is birational to $\mathbb{P}^{1}\times E$ for some smooth curve $E$. In case (1), we can take $E=\mathbb{P}^{1}$. In case (2), $E$ is isomorphic to $C$ which has bounded gonality and genus. Now we show that $E$ also has bounded gonality and genus in case (3). As $F$ is birationally bounded in this case, $F$ has bounded degree of irrationality by Lemma 2.17, hence $E$ also has bounded gonality which is bounded from above by the degree of irrationality of $F$; see [Bas17, Remark 2.1]. On the other hand, the irregularity of the surface $\mathbb{P}^{1}\times E$ is equal to the genus $g(E)$. Since $\mathbb{P}^{1}\times E$ is birationally bounded, there are only finitely many possible values for the irregularity of $\mathbb{P}^{1}\times E$, so $g(E)$ is also bounded from above. ∎
Boundedness of $\epsilon$-lc Fano varieties. Keeping the same notation as in Theorem 1.3 and assuming that $\dim X=3$, we know that $F$ is birational to $\mathbb{P}^{1}\times E$ for some smooth projective curve $E$. Notice that the assumptions in Theorem 1.3 for $\dim X=3$ is stronger than Theorem 1.1 by requiring additionally that $X$ is $\epsilon$-lc over the generic point of $Z$, that is, a general fibre of $X\to Z$ is an $\epsilon$-lc del Pezzo surface for the fixed $\epsilon>0$. It is well-known that $\epsilon$-lc del Pezzo surfaces are bounded for any fixed $\epsilon>0$ (cf. [Ale94]), and the family of del Pezzo surfaces is not bounded for $\epsilon=0$. This leads to the phenomenon that both the gonality and genus of $E$ are always bounded under the conditions of Theorem 1.3; see Corollary 1.5. However, dropping the boundedness condition, [BL21, Example 2.3] illustrates that the genus of the curve $E$ can be arbitrarily large. A general fibre of the klt Fano fibration constructed in [BL21, Example 2.3] is isomorphic to the weighted projective space $\mathbb{P}(1,1,n)$ with $n\geq 3$. The log discrepancy of $\mathbb{P}(1,1,n)$ is equal to $2/n$, hence the family of del Pezzo surfaces $\mathbb{P}(1,1,n)$ with $n\geq 3$ is not bounded; cf. [Bir21, Example 1.2]. The boundedness condition on general fibres of the klt Fano fibration is crucial in our proof of Theorem 1.3 in higher dimensions, that is, $\epsilon$-lc Fano varieties of dimension $d$ form a bounded family for any fixed positive real number $\epsilon>0$ and natural number $d$; see [Bir21, Theorem 1.1]. It is also clear from the discussion above that boundedness of complements (see [Bir19]) is another crucial ingredient of our proofs.
Acknowledgements. We thank professor Dan Abramovich for answering our numerous questions about the results in [ATW20]. We would like to thank Dingxin Zhang and Heer Zhao for very helpful discussions about logarithmic geometry. We would also like to thank Mao Sheng for his helpful comments. The first author was supported by a grant from Tsinghua University and a grant of the National Program of Overseas High Level Talent. The results and proofs of this paper were presented by the second author in a talk at the online Tsinghua Algebraic Geometry Seminar in October 2022.
2. Preliminaries
By saying a scheme, we mean a separated scheme of finite type over the fixed algebraically closed field $\mathbb{K}$ of characteristic zero. A variety is an irreducible, reduced, and quasi-projective scheme. For a scheme $X$, we denote by $X_{\text{red}}$ the maximal reduced closed subscheme of $X$ and call it the reduction of $X$.
Let $f\colon X\to Z$ be a morphism between schemes, where $Z$ is irreducible. If $X$ has a unique irreducible component $Y$ that dominates $Z$, we call $Y$ the main component of $X$. If this is the case, we usually just say that $Y$ is the main component of $X$ without mentioning that it is the unique irreducible component of $X$ dominating $Z$.
2.1. Contractions
A contraction is a projective morphism of schemes $f:X\to Y$ such that $_{*}\mathcal{O}_{X}=\mathcal{O}_{Y}; ff$ is not necessarily birational. In particular, $f$ has connected fibres. Moreover, if $X$ is normal, then $Y$ is also normal.
2.2. Divisors
Let $X$ be a scheme. By a divisor, we mean a Weil divisor on $X$, that is, a finite $\mathbb{Z}$-linear combination of codimension one irreducible and reduced closed subschemes of $X$. By a $\mathbb{Q}$-divisor (respectively, an $\mathbb{R}$-divisor), we mean a finite linear combination $\sum_{i}b_{i}B_{i}$ where every $B_{i}$ is a prime Weil divisor on $X$ and $b_{i}\in\mathbb{Q}$ (respectively, $b_{i}\in\mathbb{R}$). A $\mathbb{Q}$-divisor (respectively, an $\mathbb{R}$-divisor) is called $\mathbb{Q}$-Cartier (respectively, $\mathbb{R}$-Cartier) if it is a $\mathbb{Q}$-linear (respectively, an $\mathbb{R}$-linear) combination of Cartier divisors. Let $B_{1}$ and $B_{2}$ be two $\mathbb{R}$-divisors on $X$. We say that $B_{1}\sim B_{2}$ (respectively, $B_{1}\sim_{\mathbb{Q}}B_{2}$, respectively, $B_{1}\sim_{\mathbb{R}}B_{2}$) if $B_{1}-B_{2}$ is a $\mathbb{Z}$-linear (respectively, a $\mathbb{Q}$-linear, respectively, an $\mathbb{R}$-linear) combination of principal divisors.
Let $f\colon X\to Z$ be a morphism of schemes, and let $L$ and $M$ be $\mathbb{R}$-divisors on $X$. We say that $L\sim M$ over $Z$ (respectively, $L\sim_{\mathbb{Q}}M$ over $Z$, respectively, $L\sim_{\mathbb{R}}M$ over $Z$) if there is a Cartier (respectively, a $\mathbb{Q}$-Cartier, respectively, an $\mathbb{R}$-Cartier) divisor $N$ on $Z$ such that $L-M\sim f^{*}N$ (respectively, $L-M\sim_{\mathbb{Q}}f^{*}N$, respectively, $L-M\sim_{\mathbb{R}}f^{*}N$).
Let $f\colon X\to Z$ be a morphism of schemes and $D$ an $\mathbb{R}$-divisor on $X$. We say that $D$ is vertical/Z/Z if $f(\operatorname{Supp}D)$ does not contain any generic point of $Z$. If $D$ does not have any vertical/Z/Z irreducible components, we say that $D$ is horizontal/Z/Z.
For volume of a big $\mathbb{R}$-divisor, we refer the readers to [Laz04, §2.2.C] for the definition and more details.
2.3. Singularities of pairs
In this paper, we will use standard notions and results from minimal model program; cf. [KM98, BCHM10]. Here we collect some of the most fundamental definitions for clarification. A pair $(X,B)$ consists of a normal quasi-projective variety $X$ and an $\mathbb{R}$-divisor $B$ with coefficients in $[0,1]$ such that $K_{X}+B$ is $\mathbb{R}$-Cartier; in this case, we say that $B$ is a boundary. Let $W\to X$ be a log resolution of a pair $(X,B)$, and let $K_{W}+B_{W}$ be the pullback of $K_{X}+B$. Denote by $\mu_{D}B_{W}$ the coefficient of $B_{W}$ at a prime divisor $D$ on $W$, then the log discrepancy of $D$ with respect to $(X,B)$ is $1-\mu_{D}B_{W}$, and it is denoted by $a(D,X,B)$. We say that $(X,B)$ is lc (respectively, klt, respectively, $\epsilon$-lc) if $a(D,X,B)$ is $\geq 0$ (respectively, >0>0, respectively, $\geq\epsilon$) for every divisor $D$ on an arbitrary log resolution $W\to X$. Note that $a(D,X,B)$ can also be defined in the same way when the coefficients of $B$ do not belong to $[0,1]$ as long as $K_{X}+B$ is $\mathbb{R}$-Cartier.
Let $(X,B)$ be an lc pair. An lc place of $(X,B)$ is a prime divisor on some birational model of $X$ such that $a(D,X,B)=0$. An lc centre is the image on $X$ of an lc place.
For dlt pairs, we refer the readers to [KM98, §2.3] for the basic definitions and to [KK10] for the definition of dlt modifications (in particular, see [KK10, Theorem 3.1] for the existence of $\mathbb{Q}$-factorial dlt modifications).
For the basic definitions and properties of generalised pairs, we recommend the readers to [BZ16, §4] and [Bir19, §2]. For the definition and existence of $\mathbb{Q}$-factorial generalised dlt modifications (also called $\mathbb{Q}$-factorial generalised dlt models); see [Bir19, §2.13].
Singularities of pairs, such as lc singularities, can also be defined for demi-normal schemes, and the corresponding pairs with lc property are called slc pairs. We refer to [Kol13, Chapter 5] for more details about slc pairs.
The following result can be shown easily by taking common log resolutions and pulling back the pairs, and the proof is left to readers.
Lemma 2.4.
Let $X$ and $Y$ be normal quasi-projective varieties over a variety $Z$, and let $\phi\colon Y\dashrightarrow X$ be a birational map. Let $B$ be a $\mathbb{Q}$-divisor on $X$ such that $K_{X}+B$ is $\mathbb{Q}$-Cartier and $F$ a divisor over $X$. Define a divisor $B_{Y}$ on $Y$ by $K_{Y}+B_{Y}=\phi^{*}(K_{X}+B)$. Assume that $K_{X}+B\sim_{\mathbb{Q}}0/Z$. Then the log discrepancy $a(F,X,B)$ is equal to $a(F,Y,B_{Y})$.
2.5. Fano fibrations
Let $(X,B)$ be an lc pair over a variety $Z$. We say that $(X,B)$ is log Fano over $Z$ if $(K_{X}+B)$ is ample over $Z$. If $Z$ is a point, then $(X,B)$ is called a log Fano pair. In this case, if $B=0$, then $X$ is called a Fano variety. A morphism $(X,B)\to Z$ from an lc pair $(X,B)$ is a log Fano fibration if $(X,B)$ is log Fano over $Z$ and the underlying morphism $X\to Z$ is a contraction. A klt log Fano fibration $(X,B)\to Z$ is a log Fano fibration with $(X,B)$ klt. In this case, if $B=0$, we call $X\to Z$ a klt Fano fibration; see [BL21, §3].
2.6. Complements
We define complements as in [Sho00]; see also [Bir19, §2.18]. Let $(X,B)$ be a pair where $B$ is a boundary, and let $X\to Z$ be a contraction. Let $T=\left\lfloor{B}\right\rfloor$ and $\Delta=B-T$, and $n\in\mathbb{N}$ a natural number. An $n$-complement of $K_{X}+B$ over a point $z\in Z$ is of the form $K_{X}+B^{+}$ such that over some neighbourhood of $z$ we have the following properties:
  • $(X,B^{+})$ is lc,
  • $n(K_{X}+B^{+})\sim 0$, and
  • $nB^{+}\geq nT+\left\lfloor{(n+1)\Delta}\right\rfloor$.
In particular, $nB^{+}$ is an integral divisor. The $n$-complement is monotonic if $B^{+}\geq B$.
2.7. Couples
A couple $(X,D)$ consists of a quasi-projective variety $X$ and a reduced Weil divisor $D$ on $X$. This is more general than the definition given in [Bir19, §2.19] because we are not assuming $X$ to be normal nor projective. Also note that a couple is not necessarily a pair in the sense that we are not assuming $K_{X}+D$ to be $\mathbb{R}$-Cartier. In this paper, we often consider a couple $(X,D)$ equipped with a surjective projective morphism to a variety $Z$ in which case we often denote the couple as $(X/Z,D)$ or $(X,D)\to Z$.
We say a couple $(X/Z,D)$ is flat if both $X\to Z$ and $D\to Z$ are flat.
2.8. Strata and log smooth morphisms
We follow the conventions as in [HMX18, §2.1]. Let $(X,D)$ be a couple. The strata of $(X,D)$ are the irreducible components of the intersections
$D_{I}=\bigcap_{j\in I}D_{j}=D_{i_{1}}\cap\cdots\cap D_{i_{r}}$
of components of $D$, where $I=\{i_{1},\dots,i_{r}\}$ is a subset of the indices, including the empty intersection $X=D_{\emptyset}$. Every irreducible component in the strata of $(X,D)$ is called a stratum of $(X,D)$. If $(X,B)$ is a pair, then the strata of $(X,B)$ are the strata of the underlying couple $(X,D)$ where $\operatorname{Supp}D=\operatorname{Supp}B$.
If we are given a couple $(X,D)$ over a variety $T$, then we say that $(X,D)$ is log smooth over $T$ if $(X,D)$ has simple normal crossings and the strata of $(X,D)$ are smooth over $T$. If $(X,B)$ is a pair over a variety $U$, we say that $(X,B)$ is log smooth over $U$ if the underlying couple $(X,D)$ is log smooth over $U$ where $\operatorname{Supp}D=\operatorname{Supp}B$.
2.9. Bounded and birationally bounded families of couples
We say that a set $\mathcal{Q}$ of couples $(X,D)$ is bounded if there exist finitely many projective morphisms $V_{i}\to T_{i}$ of varieties and reduced divisors $C_{i}$ on $V_{i}$ such that for each $(X,D)\in\mathcal{Q}$, there exist an $i$, a closed point $t\in T_{i}$, and an isomorphism of couples $\phi\colon(V_{i,t},C_{i,t})\to(X,D)$, where $V_{i,t}$ and $C_{i,t}$ are the fibres over $t$ of the morphisms $V_{i}\to T_{i}$ and $C_{i}\to T_{i}$ respectively. In particular, if $D=0$ for all $(X,D)\in\mathcal{Q}$, we say the family $\mathcal{Q}$ consisting of projective varieties $X$ is bounded.
Let $\mathcal{P}$ be a set of projective varieties. We say that $\mathcal{P}$ is birationally bounded if there is a bounded family $\mathcal{Q}$ of projective varieties such that for every $X\in\mathcal{P}$ there exists a birational map $X\dashrightarrow Y$ to a member $Y$ in $\mathcal{Q}$. Note that our definition of birational boundedness is weaker than [Bir19, §2.19].
Let $\mathcal{Q}$ be a bounded family of projective couples, and let $(X,D)\in\mathcal{Q}$. When there is no confusion in the context, we usually say that $(X,D)$ belongs to a bounded family of projective couples, or just $(X,D)$ is bounded. Similarly, if $\mathcal{P}$ is a birationally bounded family of projective varieties and $X\in\mathcal{P}$, we usually say that $X$ belongs to a birationally bounded family of projective varieties, or just $X$ is birationally bounded.
We say a set $\mathcal{S}$ of projective pairs $(X,B)$ is bounded if there exist finitely many projective morphisms $Z_{i}\to U_{i}$, where $U_{i}$ is smooth, $Z_{i}$ is flat over $U_{i}$, and pairs $(Z_{i},\Sigma_{i})$, where the support of $\Sigma_{i}$ contains neither a component of a fibre nor a codimension one singular point of any fibre, such that for every $(X,B)\in\mathcal{S}$, there exist an $i$, a closed point $u\in U_{i}$, and an isomorphism of pairs $\psi\colon(Z_{i,u},\Sigma_{i,u})\to(X,B)$, where $Z_{i,u}$ and $\Sigma_{i,u}$ are the fibres over $u$ of the morphisms $Z_{i}\to U_{i}$ and $\Sigma_{i}\to U_{i}$ respectively. In particular, the coefficients of $B$ belong to a finite set; see [HMX18, page 873].
The following lemma is well-known and we omit its proof.
Lemma 2.10.
Let $\mathcal{Q}$ be a bounded set of lc projective pairs $(X,B)$. Then the set of lc centres of $(X,B)\in\mathcal{Q}$ is also bounded.
The result as follows will be used in the proof of Theorem 1.3 in §$4.$
Lemma 2.11.
Let $d$ be a natural number. Let $(X,B+M)$ be a projective generalised lc generalised pair with $\dim X=d$ such that $L=K_{X}+B+M$ is a nef and big integral Cartier divisor. Then $\left|mL\right|$ defines a birational map for some $m\in\mathbb{N}$ depending only on $d$.
Proof.
Taking a $\mathbb{Q}$-factorial dlt modification $f\colon Y\to X$ of $(X,B+M)$ (see [BZ16, Lemma 4.5]), we have that
$K_{Y}+B_{Y}+M_{Y}=f^{*}(K_{X}+B+M)$
where $B_{Y}$ is the sum of $f^{-1}_{*}B$ and the reduced $f$-exceptional divisors and $(Y,B_{Y}+M_{Y})$ is a $\mathbb{Q}$-factorial generalised dlt generalised pair. In particular, $(Y,0)$ is $\mathbb{Q}$-factorial and klt. Fix a small rational number $0<\epsilon\ll 1$. Let $g\colon W\to Y$ be a birational morphism extracting the divisors $F$ whose log discrepancies with respect to $(Y,0)$ are $\leq\epsilon$, then $W$ is $\epsilon$-lc. Write
$K_{W}+E=g^{*}K_{Y}.$
Then the coefficients of $E$ are $\geq 1-\epsilon$. Now, we can write
$(f\circ g)^{*}L-K_{W}=E+g^{*}(B_{Y}+M_{Y})$
which is clearly pseudo-effective. Since $(f\circ g)^{*}L$ is also a nef and big integral Cartier divisor, we can apply [Bir23, Theorem 1.1] to conclude. ∎
2.12. Relative degree
Let $f\colon X\to Z$ be a surjective projective morphism of quasi-projective varieties and let $A$ be a Cartier divisor on $X$. For a Weil divisor $D$ on $X$ we define the relative degree of $D$ with respect to $A$ as $\deg_{A/Z}D\coloneqq(D|_{F})\cdot(A|_{F})^{n-1}$ where $F$ is a general fibre of $f$ and $n=\dim F$. It is clear that this is a generic property, so the vertical/Z/Z components of $D$ do not contribute to the relative degree. Note that $F$ may not be irreducible.
2.13. Relatively bounded families of couples
We define relatively bounded families of couples (respectively, varieties) as in [Bir23a, §3]. Let $\mathcal{P}$ be a family of couples $(X/Z,D)$. We say $\mathcal{P}$ is relatively bounded if there exist natural numbers $d,r$ such that for each $(X/Z,D)\in\mathcal{P}$, $\dim X-\dim Z\leq d$, there is a very ample/Z/Z divisor $A$ on $X$ such that
$\deg_{A/Z}A\leq r\penalty 10000\ \penalty 10000\ \mbox{and}\penalty 10000\ \penalty 10000\ \deg_{A/Z}D\leq r,$
and $f_{*}\mathcal{O}_{X}(A)$ is locally free where $f$ denotes $X\to Z$. The local freeness condition is to ensure that $A$ embeds $X$ into the projective bundle $\mathbb{P}(f_{*}\mathcal{O}_{X}(A))$. The local freeness is automatic when $Z$ is a smooth curve and this is the case needed in this paper.
When $D=0$ for every $(X/Z,D)\in\mathcal{P}$, we then refer to $\mathcal{P}$ as a family of relatively bounded varieties.
Here we collect some useful results about relatively bounded families of projective couples; for proofs, see [Bir23a, §3].
Lemma 2.14 (cf. [Bir23a, Lemma 3.2]).
Let $W\to T$ be a projective morphism of varieties and $G$ a reduced Weil divisor on $W$. Let $\mathcal{P}$ be the set of couples $(Y/Z,E)$ satisfying the following:
  • $Z$ is a variety equipped with a morphism $Z\to T$,
  • $Y$ is an irreducible component of $Z\times_{T}W$ with reduced structure, mapping onto $Z$,
  • the image of $Y\to W$ is not contained in $\operatorname{Supp}G$, and
  • the horizontal/Z/Z part of $E$ is contained in $\operatorname{Supp}(G|_{Y})$ where $G|_{Y}$ is the divisorial part of the reduction of the closed subscheme $Y\times_{W}G$ of $Y$.
Then $\mathcal{P}$ is a relatively bounded set of couples.
Proof.
We can find effective Cartier divisors $G_{1},\dots,G_{r}$ on $W$ so that
$\operatorname{Supp}G=\bigcap_{i=1}^{r}\operatorname{Supp}G_{i}.$
Now, for any given $(Y/Z,E)$, we can apply [Bir23a, Lemma 3.2] from some $1\leq i\leq r$. ∎
As for bounded set of projective couples (cf. [Bir19, Lemma 2.21]), there is a universal family of varieties and divisors for a relatively bounded family of projective couples over smooth curves.
Lemma 2.15 ([Bir23a, Lemma 3.4, Lemma 3.5]).
Let $d$ and $r$ be natural numbers. Let $\mathcal{P}$ be the set of all couples $(X/Z,D)$ such that
  • (a)
    $(X,D)$ is a couple with $\dim X=d$,
  • (b)
    $f\colon X\to Z$ is a projective morphism onto a smooth curve,
  • (c)
    every component of $D$ is horizontal over $Z$, and
  • (d)
    AA is a very ample/Z/Z divisor on $X$ such that $\deg_{A/Z}A\leq r$ and $\deg_{A/Z}D\leq r$.
Then there exist finitely many projective morphisms ${V}_{i}\to T_{i}$ of varieties and reduced divisors $C_{i}$ on $V_{i}$ such that for each $(X/Z,D)\in\mathcal{P}$ and each closed point $z\in Z$, after shrinking $Z$ around $z$ if necessary, there are an $i$ and a morphism $Z\to T_{i}$ such that $X=Z\times_{T_{i}}V_{i}$ and $D=Z\times_{T_{i}}C_{i}$.
2.16. Irrationality
Given a projective variety $X$, we define the degree of irrationality of $X$ following [BDELU17] as
$\operatorname{irr}(X)=\min\bigg\{\delta>0\,\,\bigg|\,\,\begin{array}[]{c}\exists\text{ degree }\delta\text{ rational dominant map}\\ X\dashrightarrow\mathbb{P}^{n}\text{ with }\dim X=n\end{array}\bigg\}.$
To save space, we also say that the irrationality of $X$ is $\operatorname{irr}(X)$.
Lemma 2.17.
Let $\mathcal{Q}$ be a birationally bounded family of projective varieties. Then the irrationality $\operatorname{irr}(X)$ for $X$ in $\mathcal{Q}$ is bounded from above.
Proof.
By assumptions, we can assume that there is a single universal family $V\to T$ for $\mathcal{Q}$ as in §2.9. As irrationality is a birational invariant, we can assume that the varieties in $\mathcal{Q}$ are closed fibres of $V\to T$. Embed $V/T$ into a projective space $\mathbb{P}^{N}_{T}$ for some $N\in\mathbb{N}$. Take a general projection from $\mathbb{P}^{N}_{T}$ to some projective subspace of $\mathbb{P}^{N}_{T}$, then there exists a generically finite rational $T$-map $V\dashrightarrow\mathbb{P}^{n}_{T}$ for some $n\leq N$. Possibly taking a stratification of $\mathcal{Q}$ shows that the irrationality for $X\in\mathcal{Q}$ is bounded from above. ∎
3. Couples and toroidal geometry
3.1. Morphisms of couples
A morphism $(Z,E)\to(V,C)$ between couples is a morphism $f\colon Z\to V$ of schemes such that $f^{-1}(C)\subseteq E$.
3.2. Toric varieties and toric morphisms
An affine toric variety $X$ of dimension $d$ is an irreducible affine variety containing an algebraic torus $\mathbb{T}_{X}\cong\mathbb{G}_{m}^{d}$ as a Zariski open subset such that the action of $\mathbb{T}_{X}$ on itself extends to an algebraic action of $\mathbb{T}_{X}$ on $X$. When the affine toric variety $X$ is normal, this is equivalent to giving a pair $(N_{X},\sigma)$ where $N_{X}$ is a lattice of finite rank and $\sigma$ is a strongly convex rational polyhedral cone in $N_{X}\otimes_{\mathbb{Z}}\mathbb{R}$ such that $X=\operatorname{Spec}\mathbb{K}[\sigma^{\vee}\cap M_{X}]$ where $M_{X}$ is the dual lattice of $N_{X}$ and $\sigma^{\vee}\subset M_{X}\otimes_{\mathbb{Z}}\mathbb{R}$ is the dual cone of $\sigma$. A toric morphism of affine normal toric varieties $X$ and $Y$ is given by a linear map of lattices $\phi\colon N_{X}\to N_{Y}$ such that the image of the cone of $X$ under the extended map $\phi_{\mathbb{R}}\colon N_{X}\otimes_{\mathbb{Z}}\mathbb{R}\to N_{Y}\otimes_{\mathbb{Z}}\mathbb{R}$ is contained in the cone of $Y$. We refer to [CLS11] for the general theory of toric varieties.
If $D$ is the toric boundary of $X$, that is, $D$ is the reduction of complement of the big torus $\mathbb{T}_{X}$ of $X$, it is well-known that $(X,D)$ is lc and $K_{X}+D\sim 0$; cf. [CLS11, Theorem 8.2.3]. In this case, we say that $(X,D)$ is a toric couple.
3.3. Toroidal couples
Let $(X,D)$ be a couple. We say the couple is toroidal at a closed point $x\in X$ if there exist a normal affine toric variety $W$ and a closed point $w\in W$ such that there is a $\mathbb{K}$-algebra isomorphism
$\widehat{\mathcal{O}}_{{X},{x}}\to\widehat{\mathcal{O}}_{{W},w}$
of completions of local rings so that the ideal of $D$ is mapped to the ideal of the toric boundary divisor $C\subset W$, that is, the reduction of complement of the big torus $\mathbb{T}_{W}$ of $W$. We call $\{(W,C),w\}$ a local toric model of $\{(X,D),x\}$. We say the couple $(X,D)$ is toroidal if it is toroidal at every closed point.
In literature, the open immersion $U_{X}\coloneqq X\setminus\operatorname{Supp}D\subset X$ is called a toroidal embedding; for example, see [KKMS73, II §1]. We usually denote this toroidal embedding by $(U_{X}\subset X)$. Note that $U_{X}$ is smooth as $\mathcal{O}_{X,x}$ is regular if and only if $\widehat{\mathcal{O}}_{X,x}$ is regular. For citations, we will use the notions toroidal couples and toroidal embeddings interchangeably to be consistent with literature, for example, [AK00, KKMS73]. Moreover, if the embedding $(U_{X}\subset X)$, or equivalently the couple $(X,D)$, is clear from the context, we just say that $X$ is a toroidal variety.
Let $(X,D)$ be a toroidal couple. If every irreducible component of $D$ is normal, we call $(X,D)$ a strict toroidal couple. In this case, we say that the corresponding toroidal embedding $(U_{X}\subset X)$ is a strict toroidal embedding or a toroidal embedding without self-intersection; see [KKMS73, page 57].
Lemma 3.4 ([Bir23a, Lemma 3.8]).
Let $(X,D)$ be a toroidal couple. Then $X$ is normal and Cohen-Macaulay, $K_{X}+D$ is Cartier, and $(X,D)$ is an lc pair.
Proof.
See [Bir23a, §3]. ∎
Proposition 3.5.
Let $(X,D)$ be a toroidal couple, and let $F$ be a divisor over $X$. Then the log discrepancy $a(F,X,D)$ is a non-negative integer; in particular, if $F$ is a divisor over $X$ with $a(F,X,D)<1$, then $a(F,X,D)=0$. Moreover, if $(X,D)$ is a strict toroidal couple, and if $F$ is a divisor over $X$ with $a(F,X,D)=0$, then $F\to\operatorname{centre}_{X}F$ has irreducible and rational general closed fibres.
Proof.
If $(X,D)$ is a toroidal couple, by Lemma 3.4, $K_{X}+D$ is Cartier, and $(X,D)$ is an lc pair, so $a(F,X,D)$ is a non-negative integer. Now, we assume that $(X,D)$ is a strict toroidal couple and that $a(F,X,D)=0$.
Step 1. Shrinking $X$ around the generic point of $\operatorname{centre}_{X}F$, by [KKMS73, page 195], there exist a normal affine toric variety $W$ and an étale morphism $p\colon X\to W$ such that $X\times_{Y}C$ is equal to $D$, where $C$ is the reduction of complement of the torus of $W$.
Let $X^{\prime}\to X$ be a log resolution such that $F$ is a nonsingular divisor on $X^{\prime}$. Denote by $\eta_{F}$ the generic point of $F$. By [Kol13, Lemma 2.22], there exists a morphism $W_{n}\to W$, which is the composite of a sequence of blowing ups, such that
  • (i)
    there exists an induced rational map $f_{n}\colon X^{\prime}\dashrightarrow W_{n}$, and
  • (ii)
    the image $f_{n}(\eta_{F})$ is a codimension one regular point of $W_{n}$.
By taking a further log resolution of $W_{n}$, we can assume that $W_{n}$ is also smooth. Taking the fibre product gives an induced rational map $f_{n}^{\prime}\colon X^{\prime}\dashrightarrow X\times_{W}W_{n}$. We include a commutative diagram for convenience of readers as follows.
[…]
Since $p$ is étale, $X\times_{W}W_{n}$ is irreducible and smooth; moreover, the projection $X\times_{W}W_{n}\to X$ is birational. Then $f_{n}^{\prime}(\eta_{F})$ is also a codimension one regular point of $X\times_{W}W_{n}$, hence $f_{n}^{\prime}$ is an isomorphism near $\eta_{F}$ by [Liu06, Lemma 3.3.24]. Let $F_{n}$ be the closure of $f_{n}(\eta_{F})$. Then by [Kol13, (2.42.2)], $a(F_{n},W,C)=0$, that is, $F_{n}$ is a log canonical place of $(W,C)$.
Step 2. Let $\pi\colon W^{\prime}\to W$ be a toric log resolution of $W$, and let $\phi_{n}\colon W_{n}\dashrightarrow W^{\prime}$ be the induced birational map. If $C^{\prime}$ is the complement of the torus of $W^{\prime}$, then $K_{W^{\prime}}+C^{\prime}=\pi^{*}(K_{W}+C)$. Thus, the closure of $\phi_{n}(\eta_{F_{n}})$ in $W^{\prime}$, that is, $\operatorname{centre}_{W^{\prime}}(F_{n})$, is an lc centre of $(W^{\prime},C^{\prime})$, where $\eta_{F_{n}}$ is the generic point of $F_{n}$. Then $\operatorname{centre}_{W^{\prime}}(F_{n})$ is an irreducible component of the intersections of irreducible components of $C^{\prime}$ as $(W^{\prime},C^{\prime})$ is log smooth.
By the orbit-cone correspondence of toric varieties (see [CLS11, Theorem 3.2.6, Proposition 3.2.7]), each irreducible component of the intersections of prime divisors of $C^{\prime}$ is a disjoint union of finitely many orbits of the torus action, and every member of these irreducible components is a toric variety $V(\tau)=\overline{O(\tau)}$ for some orbit $O(\tau)$, where $\tau$ is a face of some cone in the fan defining $(W^{\prime},C^{\prime})$. Pick $V(\tau)$ so that $\operatorname{centre}_{W^{\prime}}(F_{n})=V(\tau)$. By blowing up $W^{\prime}$ along $V(\tau)$, we get a toric morphism $b\colon B_{V(\tau)}(W^{\prime})\to W^{\prime}$. Then $(B_{V(\tau)}(W^{\prime}),\lfloor b^{*}C^{\prime}\rfloor)$ is also log smooth and $\operatorname{centre}_{B_{V(\tau)}(W^{\prime})}(F_{n})$ is a log canonical centre of the pair $(B_{V(\tau)}(W^{\prime}),\lfloor b^{*}C^{\prime}\rfloor)$. We can blow up $\operatorname{centre}_{B_{V(\tau)}(W^{\prime})}(F_{n})$ further, and inductively the centre of $F_{n}$ on some blowing up $\widetilde{W}\to W^{\prime}$ will be a divisor generically isomorphic to $F_{n}$ by [Kol13, Lemma 2.22]. Since each step of the blowing ups is a toric morphism, we have $\widetilde{W}\to W^{\prime}$ is a toric morphism and $\widetilde{F}\coloneqq\operatorname{centre}_{\widetilde{W}}(F_{n})$ is a toric variety as a torus-invariant divisor on $\widetilde{W}$.
[…]
Denote by $\widetilde{\sigma}\in\widetilde{\Sigma}$ the ray (respectively, $\sigma\in\Sigma$ the cone) such that $\widetilde{F}=V(\widetilde{\sigma})=\overline{O(\widetilde{\sigma})}$ (respectively, $V=V(\sigma)=\overline{O(\sigma)}$). Denote by $N_{\sigma}$ the sublattice of $N$ generated as a subgroup by $\sigma\cap N$, and similarly, $\widetilde{N}_{\widetilde{\sigma}}$ for $\widetilde{\sigma}\in\widetilde{\Sigma}$. Set $N(\sigma)\coloneqq N/N_{\sigma}$ and $\widetilde{N}(\widetilde{\sigma})\coloneqq\widetilde{N}/\widetilde{N}_{\widetilde{\sigma}}$. Then the morphism of torus orbits
$\pi_{\sigma}^{\widetilde{\sigma}}\colon O(\widetilde{\sigma})\to O(\sigma)$
is induced by the surjective morphism of lattices
$\overline{\varphi}_{\sigma}^{\widetilde{\sigma}}\colon\widetilde{N}(\widetilde{\sigma})\to N(\sigma).$
Thus, the fibre $F^{\widetilde{\sigma}}_{\sigma}$ of $\pi_{\sigma}^{\widetilde{\sigma}}$ over a closed point in $O(\sigma)$ is isomorphic to the torus $\mathbb{T}_{\varphi^{-1}(N_{\sigma})/\widetilde{N}_{\widetilde{\sigma}}}$; see [HLY02, page 464]. Note that the fibre of $\pi^{\widetilde{\sigma}}_{\sigma}$ over a point of $O(\sigma)$ depends only on the orbit $O(\sigma)$; see [HLY02, Proposition 2.1.4].
By generic flatness, a general fibre of $\widetilde{F}\to V$ is pure dimensional of dimension $\dim\widetilde{F}-\dim V$, which is equal to $\dim O(\widetilde{\sigma})-\dim O(\sigma)$. Let $\widetilde{F}_{c}$ be a general fibre of $\widetilde{F}\to V$, where we can assume that $c\in O(\sigma)$. As $\dim\widetilde{F}_{c}\cap(\widetilde{F}\setminus O(\widetilde{\sigma}))<\dim\widetilde{F}_{c}$, there is an one-to-one correspondence between irreducible components of $\widetilde{F}_{c}$ and irreducible components of the open subset $O(\widetilde{\sigma})_{c}\cong F^{\widetilde{\sigma}}_{\sigma}\subset\widetilde{F}_{c}$. Therefore, a general fibre of $\widetilde{F}\to V$ is irreducible and rational.
Step 3. Since $F_{n}\dashrightarrow\widetilde{F}$ is a birational map of nonsingular varieties over $V$, a general fibre of $F_{n}\to V$ is also irreducible and rational. Denote by $F^{\prime}$ the closure of $f_{n}^{\prime}(\eta_{F})$, which is a prime divisor on $X\times_{W}W_{n}$; see Step 1. View $(\operatorname{centre}_{X}F)\times_{W}F_{n}$ as a closed subscheme of $X\times_{W}W_{n}$. As a general fibre of $F_{n}\to V$ is irreducible, $F^{\prime}$ is the only irreducible component of $(\operatorname{centre}_{X}F)\times_{W}F_{n}$ that dominates $\operatorname{centre}_{X}F$. Thus, a general fibre of $F^{\prime}\to\operatorname{centre}_{X}F$ is isomorphic to a general of $F_{n}\to V$. As $F\dashrightarrow F^{\prime}$ is a birational map over $\operatorname{centre}_{X}F$, by generic flatness, we can conclude that a general fibre of $F\to\operatorname{centre}_{X}F$ is irreducible and rational. ∎
3.6. Toroidal morphisms
Now let $(X,D)$ and $(Y,E)$ be couples, and let $f\colon X\to Y$ be a morphism of varieties. Let $x\in X$ be a closed point and $y=f(x)$. We say $(X,D)\to(Y,E)$ is a toroidal morphism at $x$ if there exist local toric models $\{(W,C),w\}$ and $\{(V,B),v\}$ of $\{(X,D),x\}$ and $\{(Y,E),y\}$ respectively, and a toric morphism $g\colon W\to V$ of normal affine toric varieties so that we have a commutative diagram
${\widehat{\mathcal{O}}_{{X},{x}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{\mathcal{O}}_{{W},w}}$${\widehat{\mathcal{O}}_{{Y},{y}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{\mathcal{O}}_{{V},v}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
where the vertical maps are induced by the given morphisms $f$ and $g$, and the horizontal maps are isomorphisms induced by the local toric models. We say the morphism $f\colon(X,D)\to(Y,E)$ is toroidal if it is toroidal at every closed point of $X$. Equivalently, we call the corresponding morphism $f\colon(U_{X}\subset X)\to(U_{Y}\subset Y)$ a toroidal morphism of toroidal embeddings; cf. §3.3.
By taking local toric models, it is easy to see the following result.
Lemma 3.7.
Let $f\colon(U_{X}\subset X)\to(U_{Y}\subset Y)$ be a toroidal morphism of toroidal embeddings. Then $f^{-1}(Y\setminus U_{Y})$ is contained in the toroidal boundary $X\setminus U_{X}$.
3.8. Toroidal varieties and logarithmic geometry
As there is no other applications of logarithmic geometry in this paper except in the proof of Theorem 3.11, we refer the readers to [Kat89, Kat94, GR18, Ogu18] for details about logarithmic geometry. For more details regarding the relation between toroidal geometry and logarithmic geometry, see [Qu25].
All the log-structures in this paper are defined on Zariski sites. A log-scheme $(X,\mathcal{M})$ consists of a scheme $X$ and a log-structure $\alpha\colon\mathcal{M}\to\mathcal{O}_{X}$; see [Kat94, §1]. A log-variety is a log-scheme whose underlying scheme is a variety. To avoid confusions, when saying logarithmically smooth log-morphisms, we mean logarithmically smooth morphisms in the category of log-schemes defined in the sense of [Kat89, Kat94, Ogu18]. However, when saying a log smooth morphism, we mean a log smooth morphism from a couple as in §2.8. The applications of these terminologies should be clear from the context.
Let $X$ be a variety and $\iota\colon U\hookrightarrow X$ an inclusion of an open subset. Denote by $\mathcal{M}_{U/X}$ the sheaf of monoids $\mathcal{O}_{X}\cap\iota_{*}\mathcal{O}_{X}^{*}$. We call $\mathcal{M}_{U/X}$ the compactifying log-structure associated to the embedding $(U\subset X)$; cf. [Ogu18, §III.1.6]. The log-variety $(X,\mathcal{M}_{U/X})$ is logarithmically regular (or log-regular to abbreviate) if and only if $(U\subset X)$ is a strict toroidal embedding; cf. [Kat94, (1.7)] and [Qu25].
Let $(U_{X}\subset X),(U_{Y}\subset Y)$ be strict toroidal embeddings. Let $(U_{X}\subset X)\to(U_{Y}\subset Y)$ be a morphism of embeddings, that is, a morphism $f\colon X\to Y$ of varieties such that $f(U_{X})\subseteq U_{Y}$. Then a morphism of embeddings $(U_{X}\subset X)\to(U_{Y}\subset Y)$ is a dominant toroidal morphism if and only if the induced log-morphism $(X,\mathcal{M}_{U_{X}/X})\to(Y,\mathcal{M}_{U_{Y}/Y})$ of log-varieties is logarithmically smooth in the category of log-schemes; see [Den13, Qu25].
3.9. Relative toroidalisation
First we recall a theorem first proved by Abramovich and Karu in [AK00], which says that any projective, surjective morphism of varieties can be made logarithmically smooth after birational modifications. This result is generalised to varieties over non-closed fields of characteristic zero in [ADK13].
Theorem 3.10 ([ADK13, Theorem 1.1]).
Let $f\colon X\to B$ be a projective, surjective morphism of varieties. Let $Z\subset X$ be a proper closed subscheme. Then there exists a commutative diagram
${(U_{X^{\prime}}\subset X^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{m_{X}}$$\scriptstyle{f^{\prime}}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${(U_{B^{\prime}}\subset B^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{m_{B}}$${B}$
where $X^{\prime}$ and $B^{\prime}$ are nonsingular varieties, and $m_{B}$ and $m_{X}$ are projective birational morphisms, such that
  • (1)
    the inclusions on the left are strict toroidal embeddings;
  • (2)
    $f^{\prime}$ is a projective toroidal morphism of toroidal embeddings;
  • (3)
    let $Z^{\prime}=m_{X}^{-1}(Z)$, then $Z^{\prime}$ is a simple normal crossings divisor, and $Z^{\prime}\subset X^{\prime}\setminus U_{X^{\prime}}$;
  • (4)
    the restriction of the morphism $m_{X}$ to $U_{X^{\prime}}$ is an open embedding.
We note that the method proving this theorem in [AK00, ADK13] is non-canonical and even a smooth generic fibre of $X\to B$ can be modified by it. In [ATW20], it is shown that there is a relatively canonical toroidalisation method that keeps the toroidal locus of $f$ unchanged; for more details, we refer the readers to [ATW20, §1]. In the next result, we show that a normalised base change of a dominant toroidal morphism is also toroidal. For more details and generalisation of this result to log-structures on small étale sites, we refer the readers to [Qu25].
Theorem 3.11.
Let $(U_{X}\subset X)$, $(U_{Y}\subset Y)$ and $(U_{Z}\subset Z)$ be strict toroidal embeddings. Let
$f\colon(U_{Y}\subset Y)\to(U_{Z}\subset Z)\,\text{ and }\,g\colon(U_{X}\subset X)\to(U_{Z}\subset Z)$
be morphisms of embeddings, that is, $f,g$ are morphisms of varieties such that $f(U_{Y})\subseteq U_{Z}$ and $g(U_{X})\subseteq U_{Z}$. Assume that $g$ is a dominant toroidal morphism of toroidal embeddings.
Denote by $W$ the normalisation of an irreducible component of $Y\times_{Z}X$ that dominates $Y$, and by $p\colon W\to Y$ and $q\colon W\to X$ the induced projection morphisms respectively.
${(U_{W}\subset W)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{q}$$\scriptstyle{p}$${(U_{X}\subset X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$${(U_{Y}\subset Y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${(U_{Z}\subset Z)}$
Let
$U_{W}\coloneqq p^{-1}(U_{Y})\cap q^{-1}(U_{X}).$
Then $U_{W}$ is a nonempty open subset of $W$, and $(U_{W}\subset W)$ is also a strict toroidal embedding. Moreover, the induced morphism of embeddings
$p\colon(U_{W}\subset W)\to(U_{Y}\subset Y)$
is a dominant toroidal morphism of toroidal embeddings.
Proof.
Step 1. We also denote by
$f\colon(Y,\mathcal{M}_{U_{Y}/Y})\to(Z,\mathcal{M}_{U_{Z}/Z})\,\text{ and }\,g\colon(X,\mathcal{M}_{U_{X}/X})\to(Z,\mathcal{M}_{U_{Z}/Z})$
the induced log-morphism of log-regular log-varieties, which should not lead to any confusion from the context. Denote by
$\mathbf{\Omega}\coloneqq(\Omega,\mathcal{M}_{\Omega})\,\text{ and }\,\mathbf{\Omega}^{\operatorname{fs}}\coloneqq(\Omega^{\operatorname{fs}},\mathcal{M}_{\Omega^{\operatorname{fs}}})$
the fibre product of $(Y,\mathcal{M}_{U_{Y}/Y})$ and $(X,\mathcal{M}_{U_{X}/X})$ along $(Z,\mathcal{M}_{U_{Z}/Z})$ in the category of log-schemes and fs log-schemes respectively. By [Ogu18, III.2.1.2], the underlying scheme $\Omega$ is isomorphic to $Y\times_{Z}X$, and $\mathcal{M}_{\Omega}$ is a coherent log-structure on $\Omega$.
By [GR18, Remark 12.2.36 (ii)], the morphism of schemes $\Omega^{\operatorname{fs}}\to\Omega$ is a finite surjective morphism. By [GR18, Proposition 12.3.34] and the proof of [GR18, Proposition 12.2.35],
$\mathbf{\Omega}^{\operatorname{fs}}\to\mathbf{\Omega}$
is log-étale. As logarithmic smoothness is stable under composition and base change (see [GR18, Proposition 12.3.24]), the induced log-morphism of log-schemes
$\mathbf{\Omega}^{\operatorname{fs}}\to(Y,\mathcal{M}_{U_{Y}/Y})$
is logarithmically smooth. Then by [Kat94, (8.2)], the log-scheme $\mathbf{\Omega}^{\operatorname{fs}}$ is log-regular. By [Kat94, (4.1)], $\Omega^{\operatorname{fs}}$ has normal irreducible components, hence $\Omega^{\operatorname{fs}}$ is a disjoint union of normal varieties. Thus, every irreducible component of $\Omega^{\operatorname{fs}}$ is a strict toroidal variety.
Let $U$ be the fibre product $U_{Y}\times_{U_{Z}}U_{X}$. By [Qu25, Lemma 6.2], $U$ is nonempty. Let $(U,\mathcal{M}_{U})$ be the log-scheme endowed with trivial log-structure. It is evident that $\mathcal{M}_{U}$ is the restriction of $\mathcal{M}_{\Omega}$ to $U$. Denote by $\Omega^{\operatorname{fs},*}$ the maximal open subset of $\Omega^{\operatorname{fs}}$ on which $\mathcal{M}_{\Omega^{\operatorname{fs}}}$ is trivial. Then by the construction in [GR18, Proposition 12.2.35], $U$ is also an open subscheme of $\Omega^{\operatorname{fs}}$ such that $\mathcal{M}_{U}=\mathcal{M}_{\Omega^{\operatorname{fs}}}|_{U}$, whence $U\subseteq\Omega^{\operatorname{fs},*}$. Moreover, by [GR18, Corollary 12.3.27], the scheme $U$ is nonsingular in the usual sense.
Step 2. Let $W^{\prime}$ be an irreducible component of $Y\times_{Z}X$ that dominates $Y$. By [Qu25, Lemma 6.2], the inverse image of $U_{X}$ to $W^{\prime}$ is a dense open subset of $W^{\prime}$. Let $W$ be the normalisation of $W^{\prime}$, and let $p\colon W\to Y$ and $q\colon W\to X$ be the projections. Then $p^{-1}(U_{Y})\cap q^{-1}(U_{X})$ is a nonempty dense open subset of $W$. Since $U$ is a nonsingular scheme, there is a unique irreducible component $U_{W}$ of $U$ contained in $W^{\prime}$. Then the normalisation $W\to W^{\prime}$ is an isomorphism over $U_{W}$. Thus, viewed as an open subset of $W$, $U_{W}=p^{-1}(U_{Y})\cap q^{-1}(U_{X})$. Let $W^{\prime\prime}$ be the irreducible component of $\Omega^{\operatorname{fs}}$ that contains $U_{W}$. As $U_{W}$ embeds into both $W^{\prime}$ and $W^{\prime\prime}$, the finite morphism $W^{\prime\prime}\to W^{\prime}$ is also birational. Since $\Omega^{\operatorname{fs}}$ has normal irreducible components, we see that $W^{\prime\prime}\cong W$.
Denote by $W^{*}\subseteq W$ the restriction of $\Omega^{\operatorname{fs},*}$ to $W$. By [GR18, Remark 12.2.8], $W^{*}$ is mapped into $U_{Y},U_{X}$ via $p\colon W\to Y,q\colon W\to X$. Then
$U_{W}\subseteq W^{*}\subseteq p^{-1}(U_{Y})\cap q^{-1}(U_{X})=U_{W}.$
Thus, $U_{W}=W^{*}$. Then $(U_{W}\subset W)$ is a strict toroidal embedding. Moreover, since $\mathbf{\Omega}^{\operatorname{fs}}\to(Y,\mathcal{M}_{U_{Y}/Y})$ is logarithmically smooth, we can conclude that $p\colon(U_{W}\subset W)\to(U_{Y}\subset Y)$ is a dominant toroidal morphism of toroidal embeddings. ∎
The following result is crucial in the proof of Theorem 1.3.
Theorem 3.12.
Let $\Phi\colon\mathfrak{X}\to\mathfrak{B}$ be a projective, surjective morphism of varieties with geometrically integral generic fibre. Let $\mathfrak{Z}_{\mathfrak{X}}\subset\mathfrak{X}$ and $\mathfrak{Z}_{\mathfrak{B}}\subset\mathfrak{B}$ be proper closed subsets respectively. Then there exists a toroidal morphism $\Phi^{\prime}$ such that:
  • (i)
    [Existence] There is a commutative diagram as follows:
    ${(U_{\mathfrak{X}^{\prime}}\subset\mathfrak{X}^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{m_{\mathfrak{X}}}$$\scriptstyle{\Phi^{\prime}}$${\mathfrak{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi}$${(U_{\mathfrak{B}^{\prime}}\subset\mathfrak{B}^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{m_{\mathfrak{B}}}$${\mathfrak{B}}$
    such that $m_{\mathfrak{B}}$ and $m_{\mathfrak{X}}$ are projective birational morphisms, the inclusions on the left are strict toroidal embeddings, $m_{\mathfrak{X}}^{-1}(\mathfrak{Z}_{\mathfrak{X}})$ is contained in $\mathfrak{X}^{\prime}\setminus U_{\mathfrak{X}^{\prime}}$, $m_{\mathfrak{B}}^{-1}(\mathfrak{Z}_{\mathfrak{B}})$ is contained in $\mathfrak{B}^{\prime}\setminus U_{\mathfrak{B}^{\prime}}$, and $\Phi^{\prime}$ is a surjective, projective toroidal morphism.
  • (ii)
    [Base change functoriality] There is a non-empty open subset $U_{\mathfrak{B}}\subset\mathfrak{B}\setminus\mathfrak{Z}_{\mathfrak{B}}$ satisfying the following property. Let $i\colon C\to\mathfrak{B}$ be a morphism from a smooth curve. Assume that the image $i(C)$ is not entirely contained in the closed subset $\mathfrak{B}\setminus U_{\mathfrak{B}}$. Then $\mathfrak{X}\times_{\mathfrak{B}}C$ admits a unique irreducible component dominating $C$. Denote by $Y$ the reduction of this irreducible component of $\mathfrak{X}\times_{\mathfrak{B}}C$ with projection morphism $f\colon Y\to C$. Consider the diagram arising from normalised base change as follows:
[…]
    where $C^{\prime}$ (respectively, $Y^{\prime}$) is the normalisation of the main component of $\mathfrak{B}^{\prime}\times_{\mathfrak{B}}C$ (respectively, of $\mathfrak{X}^{\prime}\times_{\mathfrak{B}^{\prime}}C^{\prime}$), and $U_{C^{\prime}}$ (respectively, $U_{Y^{\prime}}$) is the pullback of $U_{\mathfrak{B}^{\prime}}$ to $C^{\prime}$ (respectively, of $U_{\mathfrak{X}^{\prime}}$ to $Y^{\prime}$). Then $f^{\prime}$ is also a toroidal morphism of toroidal embeddings and all the properties in (i) are satisfied. In particular, $m_{C}$ and $m_{Y}$ are projective birational morphisms, and the inverse image of $\mathfrak{Z}_{\mathfrak{B}}$ in $C^{\prime}$ (respectively, of $\mathfrak{Z}_{\mathfrak{X}}$ in $Y^{\prime}$) is contained in $C^{\prime}\setminus U_{C^{\prime}}$ (respectively, in $Y^{\prime}\setminus U_{Y^{\prime}}$).
Proof.
Step 1. To include the closed subset $\mathfrak{Z}_{\mathfrak{X}}$ into the toroidal structure, we first modify $\Phi\colon\mathfrak{X}\to\mathfrak{B}$ by Theorem 3.10 as in the following diagram
${(U_{\mathfrak{X}^{\circ}}\subset\mathfrak{X}^{\circ})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{m_{\mathfrak{X}^{\circ}}}$$\scriptstyle{\Phi^{\circ}}$${\mathfrak{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi}$${(U_{\mathfrak{B}^{\circ}}\subset\mathfrak{B}^{\circ})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{m_{\mathfrak{B}^{\circ}}}$${\mathfrak{B}}$
where $m_{\mathfrak{X}^{\circ}}$ and $m_{\mathfrak{B}^{\circ}}$ are projective birational morphisms, the inclusions on the left are strict toroidal embeddings, $m_{\mathfrak{X}^{\circ}}^{-1}(\mathfrak{Z}_{\mathfrak{X}})$ is contained in the toroidal boundary $\mathfrak{X}^{\circ}\setminus U_{\mathfrak{X}^{\circ}}$, and $\Phi^{\circ}$ is a projective, surjective toroidal morphism of toroidal embeddings.
Step 2. Now we include the closed subset $\mathfrak{Z}_{\mathfrak{B}}$ into the toroidal boundary by an additional blowing up on $\mathfrak{B}^{\circ}$. Consider the closed subset $\mathfrak{Z}^{\circ}\subset\mathfrak{B}^{\circ}$ that is the union of the inverse image of $\mathfrak{Z}_{\mathfrak{B}}$ in $\mathfrak{B}^{\circ}$ and the toroidal boundary $\mathfrak{B}^{\circ}\setminus U_{\mathfrak{B}^{\circ}}$. Let $m_{\mathfrak{B}^{\prime}}\colon\mathfrak{B}^{\prime}\to\mathfrak{B}^{\circ}$ be a log resolution of $(\mathfrak{B}^{\circ},\mathfrak{Z}^{\circ})$ such that $\mathfrak{B}^{\prime}$ is smooth and the inverse image of $\mathfrak{Z}^{\circ}$ in $\mathfrak{B}^{\prime}$ is a simple normal crossing divisor. Set $U_{\mathfrak{B}^{\prime}}$ as the open complement $\mathfrak{B}^{\prime}\setminus m_{\mathfrak{B}^{\prime}}^{-1}(\mathfrak{Z}^{\circ})$. Then $(U_{\mathfrak{B}^{\prime}}\subset\mathfrak{B}^{\prime})$ is a strict toroidal embedding and
$m_{\mathfrak{B}^{\prime}}\colon(U_{\mathfrak{B}^{\prime}}\subset\mathfrak{B}^{\prime})\to(U_{\mathfrak{B}^{\circ}}\subset\mathfrak{B}^{\circ})$
is a morphism of embeddings.
Taking the normalised base change $\Phi^{\prime}$ of the toroidal morphism $\Phi^{\circ}$ (see Theorem 3.11), we have a commutative diagram
${(U_{\mathfrak{X}^{\prime}}\subset\mathfrak{X}^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{m_{\mathfrak{X}^{\prime}}\,\,\,\,}$$\scriptstyle{\Phi^{\prime}}$${(U_{\mathfrak{X}^{\circ}}\subset\mathfrak{X}^{\circ})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,m_{\mathfrak{X}^{\circ}}}$$\scriptstyle{\Phi^{\circ}}$${\mathfrak{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi}$${(U_{\mathfrak{B}^{\prime}}\subset\mathfrak{B}^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{m_{\mathfrak{B}^{\prime}}\,\,\,\,}$${(U_{\mathfrak{B}^{\circ}}\subset\mathfrak{B}^{\circ})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,m_{\mathfrak{B}^{\circ}}}$${\mathfrak{B}}$
where $m_{\mathfrak{X}^{\prime}}$ and $m_{\mathfrak{B}^{\prime}}$ are projective birational morphisms, $(U_{\mathfrak{X}^{\prime}}\subset\mathfrak{X}^{\prime})$ is also a strict toroidal embedding, the open subset $U_{\mathfrak{X}^{\prime}}\subset\mathfrak{X}^{\prime}$ is the intersection of the open subsets $(\Phi^{\prime})^{-1}(U_{\mathfrak{B}^{\prime}})$ and $m_{\mathfrak{X}^{\prime}}^{-1}(U_{\mathfrak{X}^{\circ}})$, and $\Phi^{\prime}$ is a surjective, projective toroidal morphism.
Denote by $m_{\mathfrak{X}}$ and $m_{\mathfrak{B}}$ the composite projective birational morphisms
$m_{\mathfrak{X}^{\circ}}\circ m_{\mathfrak{X}^{\prime}}\,\,\text{ and }\,\,m_{\mathfrak{B}^{\circ}}\circ m_{\mathfrak{B}^{\prime}}.$
By construction, $m_{\mathfrak{X}}^{-1}(\mathfrak{Z}_{\mathfrak{X}})$ (respectively, $m_{\mathfrak{B}}^{-1}(\mathfrak{Z}_{\mathfrak{B}})$) is contained in $\mathfrak{X}^{\prime}\setminus U_{\mathfrak{X}^{\prime}}$ (respectively, in $\mathfrak{B}^{\prime}\setminus U_{\mathfrak{B}^{\prime}}$). This proves (i).
Step 3. Let $V_{\mathfrak{B}}\subset\mathfrak{B}$ be the smallest closed subset such that
  • $V_{\mathfrak{B}}$ contains the image of the toroidal boundary $\mathfrak{B}^{\prime}\setminus U_{\mathfrak{B}^{\prime}}$,
  • $m_{\mathfrak{B}}$ is an isomorphism over $\mathfrak{B}\setminus V_{\mathfrak{B}}$, and
  • every fibre of $\Phi$ over $\mathfrak{B}\setminus V_{\mathfrak{B}}$ is geometrically integral.
It is evident that $U_{\mathfrak{B}}\coloneqq\mathfrak{B}\setminus V_{\mathfrak{B}}$ is a non-empty open subset of $\mathfrak{B}$. Moreover, by construction, $U_{\mathfrak{B}}$ is contained in $\mathfrak{B}\setminus\mathfrak{Z}_{\mathfrak{B}}$ as the closed subset $\mathfrak{Z}_{\mathfrak{B}}$ is contained in $V_{\mathfrak{B}}$. Then the base change functoriality for curves in (ii) also follows from Theorem 3.11. ∎
4. Bounding irrationality of degenerations
In this section, we prove Theorem 1.3 about the generic rational fibration structure of irreducible fibres on degenerations of klt Fano fibrations. First we recall the setup of this result. Fix $d\in\mathbb{N}$, $\epsilon>0$ and $t\in(0,1]$. Denote by $\mathcal{F}_{d,\epsilon,t}$ the set of data $(X/Z,tF)$ where
  • $Z$ is a smooth curve,
  • $f\colon X\to Z$ is a klt Fano fibration with $\dim X=d$, that is, $X$ is a klt variety of dimension $d$ and $K_{X}$ is ample/Z/Z,
  • $X$ is $\epsilon$-lc over the generic point $\eta_{Z}$ of $Z$,
  • $F$ is the reduction of an irreducible fibre of $f$ over a closed point $z\in Z$, and
  • $(X,tF)$ is log canonical.
Decreasing $t$ if necessary, we can assume that $t\in(0,1]$ is a rational number. Then by the existence of complements (see [Bir19, Theorem 1.8]), there exists an $n\in\mathbb{N}$ depending only on $d$ and $t$ such that there exists a monotonic $n$-complement $K_{X}+B^{+}$ of $K_{X}+tF$ over $z\in Z$. Up to shrinking $Z$ around $z$ if necessary, we can assume that the boundary $B^{+}$ is defined over the whole $Z$. More precisely, there exists a boundary $B^{+}$ on $X$ such that
  • $(X,B^{+})$ is log canonical,
  • $n(K_{X}+B^{+})\sim 0/Z$,
  • $tF\leq B^{+}$, and
  • $a(F,X,B^{+})<1$.
Note that the last condition on log discrepancy follows from that $(X,tF)$ is lc and $tF\leq B^{+}$.
4.1. Modification to generically smooth couples
We first show that we can modify the fibration $(X/Z,tF)$ to a relatively bounded fibration whose general fibres are smooth.
Lemma 4.2.
There is a relatively bounded family $\mathcal{P}$ of couples $(Y/Z,D)$ such that for each $f\colon X\to Z$ in $\mathcal{F}_{d,\epsilon,t}$ with an $n$-complement $K_{X}+B^{+}$ we have a couple $(Y/Z,D)\in\mathcal{P}$ admitting a birational $Z$-map $\phi\colon X\dashrightarrow Y$ such that
  • (1)
    the reduced divisor $D$ is horizontal over $Z$,
  • (2)
    $\phi$ is an isomorphism over the generic point of $Z$, and
  • (3)
    $\operatorname{Supp}B^{+}$ is mapped isomorphically to $\operatorname{Supp}D$ over the generic point of $Z$.
Proof.
Step 1. For each $f\colon X\to Z$ in $\mathcal{F}_{d,\epsilon,t}$, denote by $U_{f}$ the maximal open subset of $Z$ over which every fibre of $f$ is an $\epsilon$-lc Fano variety of dimension d1d-1. Denote by $\mathcal{F}_{d,\epsilon,t}^{\text{fb}}$ the set of all fibres of all $f\in\mathcal{F}_{d,\epsilon,t}$ over $U_{f}$. Denote by $\mathcal{F}_{d,\epsilon,t}^{\circ}$ the set of all morphisms that are the restrictions of $f\in\mathcal{F}_{d,\epsilon,t}$ over $U_{f}$. That is, every member in $\mathcal{F}_{d,\epsilon,t}^{\circ}$ is equal to $f|_{f^{-1}(U_{f})}$ for some $f\in\mathcal{F}_{d,\epsilon,t}$. By the boundedness of $\epsilon$-lc Fano varieties of dimension d1d-1 (see [Bir21, Theorem 1.1]), we can assume that $W\to B$ is the universal family of $\mathcal{F}_{d,\epsilon,t}^{\text{fb}}$ where $W$ and $B$ are quasi-projective varieties. By taking suitable stratifications on $\mathcal{F}_{d,\epsilon,t}^{\circ}$ and doing induction on $\dim B$, we conclude that there exists a fixed natural number $\ell$ depending only on $\mathcal{F}_{d,\epsilon,t}^{\circ}$ such that for every $f\in\mathcal{F}_{d,\epsilon,t}$ there exists an open subset $U_{f}^{\prime}\subset U_{f}$ over which $\ell K_{X}$ is very ample. Without loss of generality, we can assume that $U_{f}^{\prime}$ is equal to the open subset $U_{f}$. Moreover, by the boundedness of $\mathcal{F}_{d,\epsilon,t}^{\text{fb}}$, the relative degree of $\ell K_{X}$ over $U_{f}$ is bounded from above. Thus, $\mathcal{F}_{d,\epsilon,t}^{\circ}$ is relatively bounded.
[…]
$C+(nB^{+}-C)\sim-nK_{X}/Z.$
Thus, on a general fibre $X_{s}$ of $X\to Z$, we have
$\deg_{A/Z}C=(A|_{X_{s}})^{d-2}\cdot(C|_{X_{s}})\leq(A|_{X_{s}})^{d-2}\cdot((-nK_{X})|_{X_{s}})=(A|_{X_{s}})^{d-2}\cdot(-nK_{X_{s}})\leq r.$
Denote by $\mathcal{P}^{\circ}$ the family of all the couples $(f^{-1}(U_{f})/U_{f},C|_{f^{-1}(U_{f})})$. Then $\mathcal{P}^{\circ}$ is a relatively bounded set of couples. By Lemma 2.15, there are finitely many projective morphisms of varieties $V_{i}\to T_{i}$ and reduced divisors $C_{i}$ on $V_{i}$ such that each couple in $\mathcal{P}^{\circ}$ comes from a base change of some $(V_{i},C_{i})\to T_{i}$ (after shrinking the base $U_{f}$ if necessary). Then for every $f\colon X\to Z$ in $\mathcal{F}_{d,\epsilon,t}$, we can assume that there is a morphism from $U_{f}$ to some $T_{i}$ such that $f^{-1}(U_{f})$ (respectively, $C|_{f^{-1}(U_{f})}$) is equal to the fibre product $U_{f}\times_{T_{i}}V_{i}$ (respectively, $U_{f}\times_{T_{i}}C_{i}$).
Step 3. For every $T_{i}$, let $\widetilde{T}_{i}$ be a projective compactification of $T_{i}$. Embed $V_{i}\to T_{i}$ into some $\mathbb{P}^{n}\times T_{i}$; denote by $\widetilde{V}_{i}$ (respectively, by $\widetilde{C}_{i}$) the reduced schematic closure of $V_{i}$ (respectively, of $C_{i}$) in $\mathbb{P}^{n}\times\widetilde{T}_{i}$. Then $(\widetilde{V}_{i},\widetilde{C}_{i})\to\widetilde{T}_{i}$ is a projective compactification of $(V_{i},C_{i})\to T_{i}$ where $\widetilde{V}_{i}$, $\widetilde{C}_{i}$, and $\widetilde{T}_{i}$ are all projective. As $Z$ is a smooth curve, the morphism $U_{f}\to T_{i}$ extends to a morphism $Z\to\widetilde{T}_{i}$. Denote by $\mathcal{P}$ the set of all couples $(Y/Z,D)$ constructed from the base changes $Z\times_{\widetilde{T}_{i}}\widetilde{V}_{i}$ and $Z\times_{\widetilde{T}_{i}}\widetilde{C}_{i}$ for morphisms $Z\to\widetilde{T}_{i}$ as in Lemma 2.14. Then $\mathcal{P}$ is relatively bounded. Also, by construction, every $f\colon X\to Z$ in $\mathcal{F}_{d,\epsilon,t}$ admits a birational $Z$-map $\phi\colon X\dashrightarrow Y$ to some $(Y/Z,D)$ in $\mathcal{P}$; moreover, $(X/Z,\operatorname{Supp}B^{+})$ and $(Y/Z,\operatorname{Supp}D)$ have the same general fibres. ∎
Lemma 4.3.
There exists a relatively bounded family $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ of couples $(Y/Z,D)$ with $D$ horizontal/Z/Z such that
  • (1)
    $(Y,D)\to Z$ is generically log smooth (see §2.8),
  • (2)
    every $f\colon X\to Z$ in $\mathcal{F}_{d,\epsilon,t}$ admits a birational $Z$-map $\phi\colon X\dashrightarrow Y$ to some $Y\to Z$ in $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$, and
  • (3)
    over the generic point of $Z$, we have the following:
    • (i)
      $\phi$ does not contract any divisor, and
    • (ii)
      the divisor $D$ supports on the union of $\operatorname{Supp}\phi_{*}B^{+}$ and the support of all horizontal/Z/Z exceptional divisors of $\phi^{-1}$.
Proof.
Take the relatively bounded family of couples $\mathcal{P}$ in Lemma 4.2. Then there are finitely many universal projective morphisms of varieties as in Lemma 2.15. We can assume that there is only one such universal projective morphism $(V,C)\to T$ for $\mathcal{P}$ with $V$ and $T$ quasi-projective varieties and $C$ a reduced horizontal/T/T divisor on $V$. Take a log resolution $V^{\text{sm}}\to V$ of the couple $(V,C)$. Denote by $C^{\text{sm}}$ the reduced divisor supporting on the union of the birational transform of $C$ and all horizontal/T/T exceptional divisors of $V^{\text{sm}}\to V$.
By generic smoothness, there exists an open dense subset $U_{1}\subset T$ such that $(V^{\text{sm}},C^{\text{sm}})$ is log smooth over $U_{1}$. Moreover, we can assume that the images of all vertical/T/T exceptional divisors of $V^{\text{sm}}\to V$ do not intersect $U_{1}$. Shrinking $Z$ around the closed point $z\in Z$ if necessary, there is a morphism $Z\to T$ as in Lemma 2.15. If the image of $Z\to T$ is not entirely contained in $T_{1}\coloneqq T\setminus U_{1}$, we can take the couple $(Y/Z,D)$ as the reduction of the main component of $V^{\text{sm}}\times_{T}Z$ equipped with the reduced divisor which is the reduction of the horizontal/Z/Z part of $C^{\text{sm}}\times_{T}Z$. Write $T_{1}$ as a union of finitely many irreducible varieties, then the result follows by making a suitable stratification on $\mathcal{P}$ and doing induction on $\dim T$. ∎
4.4. Proof of bounded irrationality of irreducible fibres
Theorem 4.5 (=Theorem 1.3).
Fix positive real numbers $\epsilon>0$, $t\in(0,1]$ and a natural number $d$. Assume that $f\colon X\to Z$ is a klt Fano contraction with $\dim X=d$ such that
  • (1)
    $Z$ is a smooth curve,
  • (2)
    $X$ is $\epsilon$-lc over the generic point of $Z$, and
  • (3)
    $F$ is the reduction of an irreducible fibre of $f$ and $(X,tF)$ is lc.
Then there is a dominant rational map $F\dashrightarrow C$ whose general fibres are irreducible and rational and $C$ is a bounded smooth projective variety hence with bounded degree of irrationality.
Proof.
Decreasing $t$ if necessary, we can assume that $t$ is a rational number. Then we construct the $n$-complement $K_{X}+B^{+}$ and the family of projective couples $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ as in §4.1.
Step 1. Recall that the family of projective couples $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ constructed in Lemma 4.3 is relatively bounded. Denote by $(\mathfrak{X}_{i},\mathfrak{D}_{i})\to\mathfrak{B}_{i}$ the collection of universal families of couples of $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ as in Lemma 2.15. We can assume that there is only one such projective morphism
$\Phi\colon(\mathfrak{X},\mathfrak{D})\to\mathfrak{B}.$
By the construction in the proof of Lemma 4.3, there exists a maximal open subset $\mathfrak{B}^{\circ}\subset\mathfrak{B}$ such that $(\mathfrak{X},\mathfrak{D})$ is log smooth over $\mathfrak{B}^{\circ}$. Moreover, after shrinking $Z$ around the closed point $z\in Z$ if necessary, we can assume that every $(X/Z,tF)\in\mathcal{F}_{d,\epsilon,t}$ admits a morphism $Z\to\mathfrak{B}$ whose image intersects $\mathfrak{B}^{\circ}$. That is, for every $(X/Z,tF)\in\mathcal{F}_{d,\epsilon,t}$, there is a commutative diagram for some $(Y/Z,D)\in\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ as follows.
[…]
If the centre of $F$ on $\mathfrak{X}$ is an irreducible component of a fibre of $\Phi$, then $F$ is birationally bounded, hence $F$ has bounded irrationality by Lemma 2.17. In this case, we take $C$ to be a resolution of the centre of $F$ on $\mathfrak{X}$. Note that given a bounded family of varieties there is a bounded family of resolutions for such varieties, hence we can take $C$ to be a bounded smooth projective variety. So, in the rest of the proof, we can assume that the centre of $F$ on $\mathfrak{X}$ is not an irreducible component of a fibre of $\Phi$.
Assume that the image of $z\in Z$ in $\mathfrak{B}$ is a closed point of $\mathfrak{B}^{\circ}$. Shrinking $Z$ around $z$ if necessary, we can assume that the whole image of $Z$ in $\mathfrak{B}$ is contained in the open subset $\mathfrak{B}^{\circ}$. Then by the construction of $\mathfrak{B}^{\circ}$, the couple $(Y,D)$ is log smooth. Write
$K_{Y}+B_{Y}=(\phi^{-1})^{*}(K_{X}+B^{+}).$
By construction of the couple $(Y,D)$ in Lemma 4.3, we have $B_{Y}\leq D$. Thus, by Lemma 2.4,
$0\leq a(F,Y,D)\leq a(F,Y,B_{Y})=a(F,X,B^{+})<1,$
which implies that $a(F,Y,D)=0$ by Proposition 3.5; in particular, we conclude that $\operatorname{centre}_{Y}F$ is an lc centre of $(Y,D)$ contained in the fibre $Y_{z}$, so $\operatorname{centre}_{Y}F$ is a stratum of the log smooth couple $(Y_{z},D_{z})$. Then it is clear that $\operatorname{centre}_{Y}F$ is bounded. Moreover, as $(Y,D)$ is strict toroidal, we can conclude by applying Lemma 2.17 and Proposition 3.5 and by taking $C$ to be a bounded smooth resolution of $\operatorname{centre}_{Y}F$.
Denote by $\mathfrak{Z}_{\mathfrak{B}}\subset\mathfrak{B}$ the proper closed subset $\mathfrak{B}\setminus\mathfrak{B}^{\circ}$. From now on, we assume that the image of the closed point $z\in Z$ in $\mathfrak{B}$ is contained in the closed subset $\mathfrak{Z}_{\mathfrak{B}}$.
Step 2. By Theorem 3.12, there is a commutative diagram
${(U_{\mathfrak{X}^{\prime}}\subset\mathfrak{X}^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\,\,\,\,\,\,\,\,\,\,\,\,\,\,m_{\mathfrak{X}}}$$\scriptstyle{\Phi^{\prime}}$${\mathfrak{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi}$${(U_{\mathfrak{B}^{\prime}}\subset\mathfrak{B}^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\,\,\,\,\,\,\,\,\,\,\,\,\,\,m_{\mathfrak{B}}}$${\mathfrak{B}}$
such that $m_{\mathfrak{B}}$ and $m_{\mathfrak{X}}$ are projective birational morphisms, the inclusions on the left are strict toroidal embeddings, $m_{\mathfrak{X}}^{-1}(\mathfrak{D})$ is contained in $\mathfrak{X}^{\prime}\setminus U_{\mathfrak{X}^{\prime}}$, $m_{\mathfrak{B}}^{-1}(\mathfrak{Z}_{\mathfrak{B}})$ is contained in $\mathfrak{B}^{\prime}\setminus U_{\mathfrak{B}^{\prime}}$, and $\Phi^{\prime}$ is a toroidal morphism of toroidal embeddings. Moreover, without loss of generality, we can assume that the open subset $U_{\mathfrak{B}}$ in Theorem 3.12 (ii) is contained in $\mathfrak{B}^{\circ}$ of Step 1.
Let $Z\to\mathfrak{B}$ be a morphism for a couple $(Y/Z,D)$ in $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ which corresponds to some $(X/Z,tF)\in\mathcal{F}_{d,\epsilon,t}$. We can assume that both $m_{\mathfrak{X}}$ and $m_{\mathfrak{B}}$ are birational over the generic point of the image of $Z$ in $\mathfrak{B}$ and assume that the image of $Z$ intersects the open subset $U_{\mathfrak{B}}\subset\mathfrak{B}$; otherwise, we make a stratification of $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ and do induction on $\dim\mathfrak{B}$.
Denote by $g$ the morphism $Y\to Z$. Let $Z^{\prime}$ be the normalisation of the main component of $\mathfrak{B}^{\prime}\times_{\mathfrak{B}}Z$. Then $Z^{\prime}\to Z$ is an isomorphism via $m_{\mathfrak{B}}$ as $Z$ is a smooth curve. Let $Y^{\prime}$ be the normalisation of the main component of $\mathfrak{X}^{\prime}\times_{\mathfrak{B}^{\prime}}Z^{\prime}$, and let $U_{Y^{\prime}}\subset Y^{\prime}$ be the pullback of $U_{\mathfrak{X}^{\prime}}$ to $Y^{\prime}$. Similarly, we can define the open subset $U_{Z^{\prime}}\subset Z^{\prime}$. Denote by $h_{Y^{\prime}}$ the induced morphism $Y^{\prime}\to\mathfrak{X}^{\prime}$ and by $g^{\prime}$ the morphism $Y^{\prime}\to Z^{\prime}$. Then by Lemma 3.7, it is clear that $U_{Y^{\prime}}=h_{Y^{\prime}}^{-1}(U_{\mathfrak{X}^{\prime}})$ is equal to $h_{Y^{\prime}}^{-1}(U_{\mathfrak{X}^{\prime}})\cap(g^{\prime})^{-1}(U_{Z^{\prime}})$. By Theorem 3.12 (ii), the induced morphism $g^{\prime}\colon(U_{Y^{\prime}}\subset Y^{\prime})\to(U_{Z^{\prime}}\subset Z^{\prime})$ is also a toroidal morphism of strict toroidal embeddings. We include the commutative diagram here for convenience.
[…]
Recall that there is a birational $Z$-map $\phi\colon X\dashrightarrow Y$ and that $F$ is the reduction of the central fibre of $X$ over the closed point $z\in Z$. Moreover, by the construction in Lemma 4.3, $\phi$ does not contract any horizontal/Z/Z divisors, and all the horizontal/Z/Z exceptional divisors of $\phi^{-1}$ are contained in $\operatorname{Supp}D$. Furthermore, by Theorem 3.12 (ii), $m_{Y}^{-1}(\operatorname{Supp}D)$ is contained in $Y^{\prime}\setminus U_{Y^{\prime}}$.
Step 3. By Lemma 3.7 and the choice of the closed point $z\in Z$, the reduced divisor $\{z\}$ and the support of the fibre of $g^{\prime}$ over $z$ are contained in the toroidal boundaries of $(U_{Z^{\prime}}\subset Z^{\prime})$ and $(U_{Y^{\prime}}\subset Y^{\prime})$ respectively. Denote by $D^{\prime}$ the reduced divisor supporting on the complement $Y^{\prime}\setminus U_{Y^{\prime}}$. Define the divisor $D_{Y^{\prime}}$ by $K_{Y^{\prime}}+D_{Y^{\prime}}=(\phi^{-1}\circ m_{Y})^{*}(K_{X}+B^{+})$. Similarly, denote by $D_{Y}$ the divisor defined by $K_{Y}+D_{Y}=(\phi^{-1})^{*}(K_{X}+B^{+})$. Note that $D_{Y^{\prime}}$ and $D_{Y}$ can have components with negative coefficients. Since $Y$ is generically smooth over $Z$, over the generic point of $Z$, we can write
$K_{Y^{\prime}}+R=m_{Y}^{*}K_{Y}$
where $R\leq 0$ supports in the exceptional locus of $m_{Y}$, and hence we have
$D_{Y^{\prime}}=m_{Y}^{*}D_{Y}+R\leq m_{Y}^{*}D_{Y}$
over the generic point of $Z$. By the construction of the divisor $D$ on $Y$ in Lemma 4.3, $\operatorname{Supp}D_{Y}$ is contained in $\operatorname{Supp}D$ over the generic point of $Z$. Moreover, since $m_{Y}^{-1}(\operatorname{Supp}D)$ is contained in $\operatorname{Supp}D^{\prime}$ by construction, we conclude that $\operatorname{Supp}m_{Y}^{*}D_{Y}$ is contained in $\operatorname{Supp}D^{\prime}$ over the generic point of $Z$. Then as $D_{Y^{\prime}}\leq 1$, we have $D_{Y^{\prime}}\leq D^{\prime}$ over the generic point of $Z$. On the other hand, as the support of the fibre of $g^{\prime}$ over $z\in Z$ is contained in the toroidal boundary $D^{\prime}$, shrinking $Z$ around $z$ if necessary, we can assume that $D_{Y^{\prime}}\leq D^{\prime}$ over the whole $Z$.
Then by Lemma 2.4, we have the relation of log discrepancies
$0\leq a(F,Y^{\prime},D^{\prime})\leq a(F,Y^{\prime},D_{Y^{\prime}})=a(F,X,B^{+})<1.$
Since $(Y^{\prime},D^{\prime})$ is a strict toroidal couple, $K_{Y^{\prime}}+D^{\prime}$ is Cartier by Lemma 3.4, so $a(F,Y^{\prime},D^{\prime})=0$. That is, $\operatorname{centre}_{Y^{\prime}}(F)$ is an lc centre of the pair $(Y^{\prime},D^{\prime})$.
[…]
Step 5. Now let $V$ be an lc centre of $(Y^{\prime},D^{\prime})$ that is contained in $S$. We show that $V$ is birationally bounded. Notice that a general fibre of $\Phi^{\prime}$ is normal, hence, over the generic point of $Z$, the normal scheme $Y^{\prime}$ is isomorphic to the fibre product $\mathfrak{X}^{\prime}\times_{\mathfrak{B}^{\prime}}Z^{\prime}$. Thus, by the construction in Step 3, there is an integer $r\in\mathbb{N}$ such that the volume of $K_{Y^{\prime}}+D^{\prime}+A$ on a general fibre of $Y^{\prime}\to Z^{\prime}$ is less than $r$. Let $W$ be the normalisation of $V$. By [FH23, Theorem 1.1, Definition 1.3], we can write $(K_{Y^{\prime}}+D^{\prime})|_{W}=K_{W}+D_{W}+M_{W}$ where $(W,D_{W}+M_{W})$ is an NQC generalised lc generalised pair. Now let $L=K_{Y^{\prime}}+D^{\prime}+A$, and let $L_{W}=L|_{W}$ which is an ample Cartier divisor on $W$ as $L$ is ample and Cartier. By applying Lemma 2.11 and its proof to $L_{W}=K_{W}+D_{W}+M_{W}+A_{W}$, we see that $\left|mL_{W}\right|$ is birational for some $m\in\mathbb{N}$ depending only on $\dim W$. It is then enough to show that the volume of $L_{W}$ is bounded by [HMX13, Lemma 2.4.2 (2)].
Pick a component $S^{\prime}$ of $S$ containing $V$ and let $T$ be its normalisation. Applying [FH23, Theorem 1.1], we can write $(K_{Y^{\prime}}+D^{\prime})|_{T}=K_{T}+D_{T}$ where $(T,D_{T})$ is lc (since $T$ is codimension one there is no nef part of the generalised pair in this adjunction). Take a general Cartier divisor $P\geq 0$ on $Y^{\prime}$ containing $V$ and avoiding other lc centres of $(Y^{\prime},D^{\prime})$. For any small $\alpha>0$, $(Y^{\prime},D^{\prime}+\alpha P)$ is not lc near $V$, hence applying [FH23, Theorem 1.1] again we deduce that $(T,D_{T})$ has an lc centre $V_{T}$ mapping onto $V$.
Denote by $A_{T}$ the pullback of the ample Cartier divisor $A$ to $T$. By assumption on volumes, we have that the volume of the ample Cartier divisor $K_{T}+D_{T}+A_{T}=(K_{Y^{\prime}}+D^{\prime}+A)|_{T}$ is $\leq r$, so the volume of $K_{T}+D_{T}+A_{T}$ takes only finitely many values. Changing the ample Cartier divisor $A_{T}$ linearly, we can assume that $(T,D_{T}+A_{T})$ is lc. Then [HMX18, Theorem 1.1] shows that $(T,D_{T}+A_{T})$ belongs to a bounded set of pairs, so the lc centre $V_{T}$ of $(T,D_{T})$ is also bounded by Lemma 2.10. This implies that the volume of $(K_{Y^{\prime}}+D^{\prime}+A)|_{V_{T}}$ is bounded, hence the volume of $(K_{Y^{\prime}}+D^{\prime}+A)|_{W}$ is bounded as desired.
Step 6. Now we can conclude the proof as follows. Recall that $\operatorname{centre}_{Y^{\prime}}(F)$ is an lc centre of the pair $(Y^{\prime},D^{\prime})$ that is contained in $S$. By Step 5, $\operatorname{centre}_{Y^{\prime}}(F)$ is birationally bounded. Denote by $C$ a bounded nonsingular resolution of $\operatorname{centre}_{Y^{\prime}}(F)$, hence the irrationality $\operatorname{irr}(C)$ is bounded from above by Lemma 2.17. Moreover, by Proposition 3.5, $F\to\operatorname{centre}_{Y^{\prime}}(F)$ has irreducible and rational general fibres, so does the rational map $F\dashrightarrow C$. ∎
Corollary 4.6 (=Theorem 1.4).
Fix a positive real number $\epsilon>0$ and a natural number $d$. Let $f\colon X\to Z$ be a Fano fibration with $\dim X=d$ such that
  • $Z$ is a smooth curve, and
  • $X$ is $\epsilon$-lc.
Let $F$ be an irreducible component of a closed fibre of $f$. Then there is a dominant rational map $F\dashrightarrow C$ whose general fibres are irreducible and rational, and $C$ is a bounded smooth projective variety hence with bounded degree of irrationality.
Proof.
Let $z\in Z$ be the closed point so that $F$ is contained in the fibre $f^{-1}(z)$. By [BCHM10], we can run $\phi\colon X\dashrightarrow X^{\prime}$ a (F)(-F)-MMP over $Z$ that ends with a good minimal model of F-F. By negativity lemma, $F$ is not contracted by $\phi$. Let $F^{\prime}\coloneqq\phi_{*}F$. Then $F^{\prime}$ supports on the whole fibre of $X^{\prime}\to Z$ over $z$. Note that $K_{X^{\prime}}$ is ample over the generic point of $Z$, so $K_{X^{\prime}}$ is big/Z/Z. Again by [BCHM10], we can run $\varphi\colon X^{\prime}\dashrightarrow X^{\prime\prime}$ a $(-K_{X^{\prime}})$-MMP over $Z$ that ends with a Fano fibration $g\colon X^{\prime\prime}\to Z$. As $F^{\prime}$ supports on the whole fibre of $X^{\prime}\to Z$ over $z$, we see that $F^{\prime\prime}\coloneqq\varphi_{*}F^{\prime}$ is an irreducible divisor supporting on the whole fibre of $g$ over $z$. Up to decreasing $\epsilon$ slightly, we can take an effective $\mathbb{Q}$-divisor $B$ so that $(X,B)$ is $\epsilon$-lc, and $K_{X}+B\sim_{\mathbb{Q}}0/Z$. Let $B^{\prime\prime}$ be the pushdown of $B$ to $X^{\prime\prime}$. As $\phi,\varphi$ are birational contractions, $(X^{\prime\prime},B^{\prime\prime})$ is also $\epsilon$-lc, and $K_{X^{\prime\prime}}+B^{\prime\prime}\sim_{\mathbb{Q}}0/Z$. By [Bir23a, Theorem 1.1], the lc threshold of $g^{*}z$ with respect to $(X^{\prime\prime},B^{\prime\prime})$ is bounded from below. Then there exists a $t>0$ depending only on $d,\epsilon$ such that $(X^{\prime\prime},B^{\prime\prime}+tF^{\prime\prime})$ is lc, hence $(X^{\prime\prime},tF^{\prime\prime})$ is lc. Thus, we can conclude by Theorem 1.3. ∎